OpenStudy (anonymous):

solve: log21 (x-84) = 3- log21 (x)

ganeshie8 (ganeshie8):

collect all logs on one side

ganeshie8 (ganeshie8):

and change the equation to exponent form

ganeshie8 (ganeshie8):

start by adding log21 (x) both sides

OpenStudy (anonymous):

x-84 = 9261 - (x) ??

ganeshie8 (ganeshie8):

Very good try, but no !

ganeshie8 (ganeshie8):

you're very close :) could u show how u got that ha ?

OpenStudy (anonymous):

YEAH... but i think i got something better... u said to move all logs to one said.. so i got: log21 (x^2-84x) = 3

OpenStudy (anonymous):

because I know that what I did its wrong.. cause I get stuck there with a wrong answer..

ganeshie8 (ganeshie8):

looks better :)

OpenStudy (anonymous):

then I will eliminate the log21 and get a quadratic eq right? x^2-84x-9261 = 0 and just solve for it i guess

ganeshie8 (ganeshie8):

Excellent ! and you know very well how to solve a quadratic

OpenStudy (anonymous):

yah :) ... let me do it quick

ganeshie8 (ganeshie8):

okay, whenever u solve log/radical equations u may get `extraneous solutions` so you may need to choose the right one...

OpenStudy (anonymous):

yeha... solving it i got 148.8 , but my answer on the book says 147 ?.?

ganeshie8 (ganeshie8):

for example : log cannot suck in negative numbers, so u need to make sure that u dont input negative numbers into log

OpenStudy (anonymous):

that site is niceeee... yeah i got that... perfect.. jsut misscalculated.. :D thxs once again :)

ganeshie8 (ganeshie8):

np.. u wlc :)